OpenStudy (anonymous):

The sum of two numbers is 57 and the difference is 9. What are the numbers? (There are two, btw).

OpenStudy (anonymous):

Can you guide me? I have filled out a whole page.

hero (hero):

If x is one of the numbers and y is the other number, then: x + y = 57 x - y = 9 Solve the system.

OpenStudy (anonymous):

x+y=57 x-y=9 add to find x and then y

OpenStudy (anonymous):

Oh, okay that sounds pretty simple. I'll vote. Thank you. I haven't done this in a while.